# Estimating Semantic Alphabet Size for LLM Uncertainty Quantification

### v1

@@ 6.1 Conclusion
Several approaches for estimating LLM uncertainty have emerged in recent years, but UQ performance can be constrained absent white-box LLM access. Furthermore, the practicality of sampling-based UQ methods is limited by the computational costs associated with repeated text generation. We illustrate the importance of semantic alphabet size in LLM uncertainty estimation with small sample sizes. Adjusting discrete semantic entropy for sample coverage using our proposed semantic alphabet size estimator results in more accurate semantic entropy estimation and improved LLM incorrectness classification, compared to other discrete semantic entropy estimators. Further, the aforementioned semantic alphabet size estimator achieves as good or better performance on incorrectness classification than all nine other black-box UQ methods considered in our QA experiments.
@@ 6.2 Limitations
We highlight several limitations in the present work. First, computational constraints limit our ability to run the experiments herein with models larger than 9 billion parameters. That said, our analysis is extensive, spanning four distinct model families, four datasets, and ten uncertainty estimators. Though our results are empirical, and experimental results may vary across studies, we make efforts to express our statistical uncertainty, and our final rank-generating procedure aims to account for it.
Second, our work aims to improve upon plugin discrete semantic entropy, whose estimator is negatively biased. For this reason, we take semantic cluster labels as fixed, without ablating across alternative clustering strategies. End-to-end semantic entropy calculation involves several steps, however, and a biased estimator does not necessitate that the entire process results in a negatively biased estimate. For instance, a high false negative rate in the assignment of responses to semantic equivalence classes (e.g., due to minor lexical alterations inducing false negatives) may positively bias the final estimate (Grewal et al., 2024).
@@
Acknowledgments
@@ Appendix A Additional implementation details
A.1 Text generation hyperparameters
With the exception of Figure 2, which iterates over several sample sizes, the results shown in the main body of this work use a sample size of $n=10$, which is also found in prior work on semantic entropy (Kuhn et al., 2023; Farquhar et al., 2024; Nikitin et al., 2024; Nguyen et al., 2025). Due to its reduced computational burden compared to larger sample sizes and prevalence in the relevant literature, we will refer to $n=10$ as a “practical” or “typical” sample size. Throughout this work, we generate a maximum of 100 tokens for each response, with early stopping enabled if the EOS token is generated. All otherwise unspecified hyperparameters are set to their defaults (e.g., we do not invoke top-$k$ sampling, top-$p$ sampling, beam search, repetition penalty, etc.).
For incorrectness classification, we follow prior work that elicits concise responses by pre-pending QA queries with the following “pre-prompt” (Farquhar et al., 2024; Nikitin et al., 2024):
@@ A.2 Synthetic data generation
Above, [NUM_ANSWERS] is an integer between 1 and 50, and the available function is a Pydantic object enforcing structured generation (Colvin et al., 2024). A human annotator assessed the number of semantic categories for possible correct answers, discarded those with only one possible correct semantic category, and removed duplicate questions. The resulting dataset has $131$ unique questions. We generate $100$ responses to each question from the POTATO dataset using GPT-4o-mini. A human annotator manually assigned the responses to semantic categories.
@@ A.3 Bidirectional entailment clustering
$|S|>1$ $|S|\geq 1$
Classifier FMI NMI PA FMI NMI PA
LLM 0.84 0.72 0.84 0.86 0.63 0.83
NLI-NQ 0.84 0.64 0.85 0.85 0.63 0.83
Table 2: Evaluating semantic classification methods against human-annotated ground truth. Results are averaged across questions from the POTATO dataset, either excluding or including questions for which all responses were semantically equivalent ($|S|>1$ and $|S|\geq 1$, respectively), according to the human rater. Metrics considered are the Fowlkes–Mallows index (FMI), normalized mutual information (NMI), and pairwise agreement (PA), assessed between the semantic equivalence classes assigned by the provided method and the human-annotated ground truth. Values are rounded to two decimal places. The best-performing BEC method by each metric (unrounded) is in bold. Instances where a Semantic Embedding method outperforms a BEC method are underlined. Responses are generated by GPT-4o-mini.
@@
The BEC approach of Kuhn et al. (2023) classifies $(d_{i},d_{j})$ using a DeBERTa model fine-tuned for NLI tasks. [^2] https://huggingface.co/microsoft/deberta-large-mnli Our implementation, used throughout this work, updates the classifier to a fine-tuned NLI model based on Microsoft’s newer DeBERTaV3 (He et al., 2021). [^3] https://huggingface.co/cross-encoder/nli-deberta-v3-base The NLI classification can be performed with or without the source questions (NLI-Q and NLI-NQ, respectively); in the former, the query is prepended to both $d_{i}$ and $d_{j}$ before passing through the NLI model.
@@
To compare the alignment between each clustering method and the human-annotated ground-truth, we measure the Fowlkes–Mallows index (FMI) (Fowlkes & Mallows, 1983) and normalized mutual information (NMI) (Danon et al., 2005; Lancichinetti et al., 2009). We also consider pairwise agreement (PA), wherein we iterate over pairs of passages, classify each pair as “entailment” or “contradiction,” based on the clustering results, and report for each method the portion of pairs whose entailment label agrees with that of the human annotator (Farquhar et al., 2024). Across metrics, we observe that the NLI method without the inclusion of questions performs roughly similarly to LLM (Table 2). For its comparable performance while limiting costs, we use NLI-NQ throughout this work.
@@ A.5 Uncertainty estimators
The so-called “total uncertainty” associated with a prediction is the information entropy of the corresponding predictive posterior distribution (Malinin & Gales, 2021). In UQ for natural language generation, an analogue is predictive entropy (PE), the entropy of the distribution of generable sequences by an LLM in response to a prompt (Kuhn et al., 2023; Lin et al., 2024; Farquhar et al., 2024). Since not all generable sequences are typically known, PE is calculated using the above plugin method to estimate the entropy of a prompt-model pair’s “answer distribution” (Kadavath et al., 2022). A shortcoming of this approach is that it treats all lexically distinct sequences as unique elements of an alphabet, even if they are semantically similar.
$a$
Method 0 $0.01$ $0.1$ 11
$\widehat{|S|}_{Hybrid}$ [1, 2] [1, 2] [1, 2] [1, 2]
NumSets [7, 10] [7, 10] [7, 10] [7, 10]
Good-Turing [7, 10] [7, 10] [7, 10] [7, 10]
$U_{EigV}$ [3, 4] [3, 4] [3, 4] [3, 4]
$\widehat{\mathbb{H}}_{Plugin}$ [7, 10] [7, 10] [7, 10] [7, 10]
$\widehat{\mathbb{H}}_{CS}$ [7, 10] [7, 10] [7, 10] [7, 10]
$\widehat{\mathbb{H}}_{Hybrid}$ [4, 5] [4, 5] [4, 5] [4, 5]
PE [3, 5] [3, 5] [3, 5] [3, 5]
SNNE [6, 6] [6, 6] [6, 6] [6, 6]
KLE [1, 2] [1, 2] [1, 2] [1, 2]
@@
Although the concept of von Neumann entropy has its roots in quantum physics (Petz, 2001), it has also been used more recently in network science literature to quantify graph complexity (Minello et al., 2019).
Nikitin et al. (2024) report the heat kernel as best-performing overall:
@@
Our implementation of KLE invokes $K_{heat}$ with hyperparameter =0.3=0.3, which the authors consider a “reasonable default” that outperforms prior approaches without additional hyperparameter optimization (Nikitin et al., 2024).
@@ A.6 Bradley-Terry confidence intervals
Gao et al. (2023) offers a procedure for ascertaining CIs about ranks of Bradley-Terry strength estimates, which we cursorily summarize as follows: Assuming a Bradley-Terry model, let $\beta=\{\beta_{1},\dots,\beta_{m}\}$ be the true latent strengths of $m$ methods, and let $\widehat{\beta}=\{\widehat{\beta}_{1},\dots,\widehat{\beta}_{m}\}$ be MLEs of the same. Given a provided level $\alpha$ and a “target” method $i$, we may establish a $(1-\alpha)$ CI about $\widehat{\beta}_{i}$, as well as “slightly more conservative” intervals about the remaining strength estimates. Let $n_{1}$ be the number of resultant intervals whose lower bounds are greater than the upper bound of method $i$’s. Similarly, let $n_{2}$ be the number of intervals whose upper bounds are less than the lower bound of method $i$’s. The true rank of method $i$’s Bradley-Terry strength is in the integer interval $n_{1}+1,n-n_{2}]$ with approximate probability $1-\alpha$. Because Proposition 4.1 of Gao et al. (2023) only holds for fixed number of methods/agents/players, we illustrate the “conservative” CIs in Figure 4C. We repeat this procedure for each UQ method to establish CIs for all strength ranks. We refer interested readers to Gao et al. (2023) for further details.
The algorithm we use to estimate strength scores employs a regularization parameter, which we take as $a=0.1$ in the main body of the paper. Table 3, which repeats the analysis of Figure 4C for varying $a$, indicates that the rank order of our Bradley-Terry results is not sensitive to this choice.
@@ Appendix B The under-sampled regime
Appendix C Additional underestimation results
Detailed results for SE underestimation, broken out across four datasets, are shown in Figure 6. Results are qualitatively similar to those in Figure 2, with $\hat{\mathbb{H}}_{Hybrid}$ consistently underestimating $SE^{*}$ less than plugin DSE for practical sample sizes.
@@ Appendix C Additional underestimation results
Figure 6: Ratio of semantic entropy estimator with varying sample size ($n=5,10,25,50,75,100$) to white-box semantic entropy with $n=100$ (denoted $SE^{*}$), with values below 11 suggesting underestimation (dotted grey line). Estimators shown are the plugin estimator of canonical discrete semantic entropy (dotted rose, indigo, green, and cyan lines) and the presented semantic entropy estimator of Equation 10 (solid rose, indigo, green, and cyan lines). Results are averaged across queries for each dataset. Log scale is used on the x-axis to highlight differences between estimators with smaller sample sizes. Instances with only one observed semantic category at $n=100$, resulting in a denominator of 0, are ignored.
Appendix D Additional incorrectness detection results

### v2

@@ 5 Related work
Figure 4: Heatmap illustrating the proportion of model-dataset pairs, rounded to the nearest integer, for which a row’s method achieved a larger AUROC point estimate than a column’s method. Uncertainty methods are organized into three groups: (i) semantic alphabet size estimators, (ii) black-box discrete semantic entropy (DSE) estimators, and (iii) other uncertainty estimators, which include white-box (PE, SE) and black-box methods (SNNE, KLE). The hybrid discrete semantic entropy (DSE) estimator consistently outperforms other explicit SE estimators.
@@ 6.1 Conclusion
Several approaches for estimating LLM uncertainty have emerged in recent years, but UQ performance can be constrained, absent white-box LLM access. Furthermore, the practicality of sampling-based UQ methods is limited by the computational costs associated with repeated text generation. We illustrate the importance of semantic alphabet size in LLM uncertainty estimation with small sample sizes. Adjusting DSE for sample coverage using our proposed semantic alphabet size estimator results in more accurate SE estimation and improved LLM incorrectness classification, compared to other explicit estimators of SE. Further, we find that semantic alphabet size estimators $U_{EigV}$ and $\widehat{|S|}_{Hybrid}$ achieve higher latent strength estimates than most UQ methods considered in our incorrectness detection experiments, with $U_{EigV}$ sharing the top rank CI with KLE, a conceptually complex approach that may be more difficult to interpret.
The overall performance of SE and plugin DSE on the incorrectness task are qualitatively similar (Figure 3), and related findings appear in prior work, such as Farquhar et al. (2024) and Nguyen et al. (2025). In fact, our experiments indicate that white-box SE tends to track plugin DSE quite closely, on average, as seen in Figure 9. We reflect on this phenomenon in Appendix Section E, where we observe nearly identical distributions induced by RBMCI and empirical class frequencies, on average. A notable difference is that SE appears to be more robust to instances where exceedingly rare semantic categories are overrepresented in the sample. The sequence log-probabilities available to SE may be otherwise underutilized, however, since SE exhibits the same underestimation pattern as plugin DSE. Future work may extend the approach described herein to the white-box setting by directly adjusting SE for unobserved semantic categories.
@@ 6.2 Limitations
We highlight several limitations in the present work. First, computational constraints limit our ability to run the experiments herein with models larger than 12 billion parameters. That said, our analysis is extensive, spanning five models, four datasets, and eleven uncertainty estimators. Though our results are empirical, and experimental results may vary across studies, we make efforts to express our statistical uncertainty, and our final rank-generating procedure aims to account for it.
Second, our work aims to improve upon plugin DSE, whose estimator is negatively biased. For this reason, we take semantic cluster labels as fixed, without ablating across alternative clustering strategies. End-to-end SE calculation involves several steps, however, and a biased estimator does not necessitate that the entire process results in a negatively biased estimate. For instance, a high false negative rate in the assignment of responses to semantic categories (e.g., due to minor lexical alterations inducing false negatives) may positively bias the final estimate (Grewal et al., 2024).
@@
6.3 Reproducibility
We outline our major experimental settings in Section 4.1, as well as evaluation metrics and baseline methods in Section 4.2. In Appendix A, we elaborate extensively on additional implementation details needed to reproduce our results, including algorithms, models, and prompt templates. Interested readers will also find a brief overview of implementation variations found in other work in Section H.
6.4 LLM Usage
@@ Appendix A Additional implementation details
A.1 Text generation
Throughout this work, we generate a maximum of 100 tokens for each response, with early stopping enabled if the EOS token is generated. All otherwise unspecified hyperparameters are set to their defaults (e.g., we do not invoke top-$k$ sampling, top-$p$ sampling, beam search, repetition penalty, etc.). For incorrectness classification, we follow prior work that elicits concise responses by pre-pending QA queries with the following “pre-prompt” (Farquhar et al., 2024; Nikitin et al., 2024):
@@ A.2 Synthetic data generation
Above, [NUM_ANSWERS] is an integer between 1 and 50, and the available function is a Pydantic object enforcing structured generation (Colvin et al., 2024). A human annotator assessed the number of semantic categories for possible correct answers, discarded those with only one possible correct semantic category, and removed duplicate questions. The resulting dataset has $131$ unique questions. We also generate $100$ responses to each question from the POTATO dataset using GPT-4o-mini. A human annotator manually assigned the responses to semantic categories.
@@ A.3 Bidirectional entailment clustering
$|S|>1$ $|S|=1$
Method Variant FMI NMI PA FMI NMI PA
Bidirectional Entailment LLM 0.84 0.72 0.84 0.86 0.71 0.80
NLI-NQ 0.84 0.64 0.85 0.89 0.61 0.89
Semantic Embedding HDBSCAN 0.68 0.52 0.69 0.70 0.41 0.60
DBSCAN ($\epsilon=0.05)$ 0.86 0.71 0.82 0.86 0.65 0.77
DBSCAN ($\epsilon=0.075)$ 0.87 0.68 0.82 0.88 0.60 0.79
DBSCAN ($\epsilon=0.1)$ 0.85 0.57 0.79 0.88 0.51 0.80
DBSCAN ($\epsilon=0.15)$ 0.79 0.27 0.66 0.87 0.25 0.76
Table 2: Evaluating semantic clustering methods against human-annotated ground truth. Results are averaged across questions from the POTATO dataset, either excluding questions for which all responses were semantically equivalent or questions eliciting multiple semantic categories ($|S|>1$ and $|S|=1$, respectively), according to the human rater. Metrics considered are the Fowlkes–Mallows index (FMI), normalized mutual information (NMI), and pairwise agreement (PA), assessed between the semantic equivalence classes assigned by the provided method and the human-annotated ground truth. Values are rounded to two decimal places. The best-performing method by each metric is in bold. Responses are generated by GPT-4o-mini.
@@
The BEC approach of Kuhn et al. (2023) classifies $(d_{i},d_{j})$ using a DeBERTa model fine-tuned for NLI tasks. [^2] https://huggingface.co/microsoft/deberta-large-mnli Our implementation, used throughout this work, updates the classifier to a fine-tuned NLI model based on Microsoft’s DeBERTaV3 (He et al., 2021). [^3] https://huggingface.co/cross-encoder/nli-deberta-v3-base The NLI classification can be performed with or without the source questions (NLI-Q and NLI-NQ, respectively); in the former, the query is prepended to both $d_{i}$ and $d_{j}$ before passing through the NLI model.
@@
To compare the alignment between each clustering method and the human-annotated ground-truth, we measure the Fowlkes–Mallows index (FMI) (Fowlkes & Mallows, 1983) and normalized mutual information (NMI) (Danon et al., 2005; Lancichinetti et al., 2009). We also consider pairwise agreement (PA), wherein we iterate over pairs of passages, classify each pair as “entailment” or “contradiction,” based on the clustering results, and report for each method the portion of pairs whose entailment label agrees with that of the human annotator (similar to the evaluation done by Farquhar et al. (2024)). For some queries, the human annotations assign all responses to the same semantic equivalence class (i.e., $|S|=1$). To ensure clustering methods are performant in both $|S|=1$ and $|S|>1$ scenarios, we calculate metrics for each (Table 2). Although this paper inherits its BEC semantic clustering strategy from prior work (Kuhn et al., 2023; Farquhar et al., 2024), we also consider several comparative baselines based on semantic embedding distance. In particular, we embed responses using bge-base-en-v1.5444https://huggingface.co/BAAI/bge-base-en-v1.5 and cluster according to their cosine distances via either density-based spatial clustering of applications with noise (DBSCAN) (Ester et al., 1996) with radius parameter $\epsilon=0.05,0.075,0.1,0.15$ or hierarchical DBSCAN (HDBSCAN) (Campello et al., 2013).
By most metrics, the best alignment with the human labels in each scenario (i.e., $|S|=1$ vs. $|S|>1$) is achieved by a BEC-based strategy (Table 2). The baselines using semantic embeddings have the additional limitation of relying on semantic similarity as a proxy for semantic equivalence. Among BEC strategies, the NLI variant without the inclusion of questions performs roughly similarly to LLM-based alternative. For its comparable performance while limiting costs, we use NLI-NQ throughout this work.
@@ A.5 Other uncertainty estimators
The so-called “total uncertainty” associated with a prediction is the information entropy of the corresponding predictive posterior distribution (Malinin & Gales, 2021). In UQ for natural language generation, an analogue is predictive entropy (PE), the entropy of the distribution of generable sequences by an LLM in response to a prompt (Kuhn et al., 2023; Lin et al., 2024; Farquhar et al., 2024). Since not all generable sequences are typically known, PE is calculated using the plugin method to estimate the entropy of a prompt-model pair’s “answer distribution” (Kadavath et al., 2022). A shortcoming of this approach is that it treats all lexically distinct sequences as unique elements of an alphabet, even if they are semantically similar.
@@
Although the concept of von Neumann entropy has its roots in quantum physics (Petz, 2001), it has also been used more recently in network science literature to quantify graph complexity (Minello et al., 2019). Nikitin et al. (2024) report the heat kernel as best-performing overall:
@@
Our implementation of KLE invokes $K_{heat}$ with hyperparameter $t=0.3$, which the authors consider a “reasonable default” that outperforms prior approaches without additional hyperparameter optimization (Nikitin et al., 2024).
Explicit estimation of semantic entropy.
We do not include SNNE or KLE, as implemented, among our list of “explicit” estimators of SE, because they do not explicitly estimate the entropy of the “true” distribution over semantic equivalence classes. For instance, Nguyen et al. (2025) describe a formulation of SNNE that equals DSE and Nikitin et al. (2024) show that a kernel exists to recover SE from KLE, but these are not necessarily the configurations used experimentally.
@@ A.6 Bradley-Terry confidence intervals
Gao et al. (2023) offers a procedure for ascertaining CIs about ranks of Bradley-Terry strength estimates, which we cursorily summarize as follows:
$a$
0 $0.01$ $0.1$ 11
$\widehat{|S|}_{Hybrid}$ [1, 4] [1, 4] [1, 4] [1, 4]
NumSets [7, 11] [7, 11] [7, 11] [7, 11]
Good-Turing [7, 11] [7, 11] [7, 11] [7, 11]
$U_{EigV}$ [1, 3] [1, 3] [1, 3] [1, 3]
$\widehat{\mathbb{H}}_{Hybrid}$ [4, 5] [4, 5] [4, 5] [4, 5]
$\widehat{\mathbb{H}}_{Plugin}$ [7, 11] [7, 11] [7, 11] [7, 11]
$\widehat{\mathbb{H}}_{CS}$ [7, 11] [7, 11] [7, 11] [7, 11]
PE [3, 4] [3, 4] [3, 4] [3, 4]
SNNE [6, 6] [6, 6] [6, 6] [6, 6]
KLE [1, 3] [1, 3] [1, 3] [1, 3]
SE [7, 11] [7, 11] [7, 11] [7, 11]
Assuming a Bradley-Terry model, let $\beta=\{\beta_{1},\dots,\beta_{m}\}$ be the true latent strengths of $m$ methods, and let $\widehat{\beta}=\{\widehat{\beta}_{1},\dots,\widehat{\beta}_{m}\}$ be MLEs of the same. Given a provided level $\alpha$ and a “target” method $i$, we may establish a $(1-\alpha)$ CI about $\widehat{\beta}_{i}$, as well as “slightly more conservative” intervals about the remaining strength estimates. Let $n_{1}$ be the number of resultant intervals whose lower bounds are greater than the upper bound of method $i$’s. Similarly, let $n_{2}$ be the number of intervals whose upper bounds are less than the lower bound of method $i$’s. The true rank of method $i$’s Bradley-Terry strength is in the integer interval $n_{1}+1,n-n_{2}]$ with approximate probability $1-\alpha$.
Because Proposition 4.1 of Gao et al. (2023) only holds for fixed number of methods/agents/players, we illustrate the “conservative” CIs in Figure 3C. We repeat this procedure for each UQ method to establish CIs for all strength ranks. We refer interested readers to Gao et al. (2023) for further details. The algorithm we use to estimate strength scores employs a regularization parameter, which we take as $a=0.1$ in the main body of the paper. Table 3, which repeats the analysis of Figure 3C for varying $a$, indicates that the rank order of our Bradley-Terry results is not sensitive to this choice.
@@ Appendix B The under-sampled regime
Appendix C Sample size and information gain
Because the true distribution over semantic equivalence classes is not known, we assume that $SE^{*}$ using $n=100$ is an adequate representation of the true estimand. In support of this choice, we consider the information gain per denary (i.e., group of ten samples, so the first denary is the first ten samples, etc.) with respect to the distribution over semantic categories - that is, the marginal information about this distribution obtained by observing each successive ten samples (Figure 6). Across models and datasets, this quantity plateaus to low values well before $n=100$. Predictably, the first denary of samples tends to provide the most information, especially for datasets whose queries can have multiple possible correct semantic categories.
Figure 6: Information gain with respect to the distribution over semantic categories, per denary (group of $10$ samples). Because the first denary is novel, its information gain is equal to its plugin entropy. Results are averaged across queries for each dataset. Log scale is used on the x-axis to highlight differences for early denaries.
Appendix D Additional underestimation results
@@
Figure 7: Ratio of semantic entropy (SE) estimator with varying sample size ($n=5,10,25,50,75,100$) to white-box SE with $n=100$ (denoted $SE^{*}$), with values below 11 suggesting underestimation (dotted grey line). Estimators shown are the plugin estimator of canonical discrete semantic entropy (DSE; dotted rose, indigo, green, and cyan lines) and the presented DSE estimator of Equation 10 (solid rose, indigo, green, and cyan lines). Results are averaged across queries for each dataset. Log scale is used on the x-axis to highlight differences between estimators with smaller sample sizes. Instances with a denominator of 0 are ignored.
Detailed results for SE underestimation, broken out across four datasets, are shown in Figure 7. Results are qualitatively similar to those in Figure 2, with $\hat{\mathbb{H}}_{Hybrid}$ typically underestimating $SE^{*}$ less than plugin DSE for practical sample sizes. Additionally, we present a version of Figure 2 that includes white-box SE (Figure 8) and elaborate on the close tracking between SE and plugin DSE in Section E.
Figure 8: Illustrating underestimation in discrete semantic entropy (DSE) calculation with typical sample sizes. The ratios of semantic entropy (SE) estimators with varying sample size ($n=5,10,25,50,75,100$) to white-box SE with $n=100$ (denoted $SE^{*}$) are shown, with values below 11 suggesting underestimation (dotted grey line). The estimators displayed are the plugin estimator of canonical DSE (i.e., Equation 5, dotted green line), the “hybrid” SE estimator of Equation 10 (solid indigo line), and white-box SE (Equation 3, dotted cyan line). Results are averaged across queries within each dataset, then uniformly averaged across datasets. Log scale is used on the x-axis to highlight differences between estimators with smaller sample sizes. Instances with a denominator of 0 are ignored.
$SE^{*}$ can be exactly zero (e.g., when only one observed semantic category is observed at $n=100$), hence we exclude such instances from Figures 2, 7, and 8. Occasionally, exceedingly rare semantic categories are revealed early, resulting in coverage-adjusted estimates that substantially exaggerate the distribution’s tail mass. We observe a few instances of this phenomenon in our sampling from the Gemma-3-12B model on the POTATO dataset: three queries in POTATO for which Gemma-3-12B elicits a non-zero $SE^{*}$ (calculated using $n=100$), but a hybrid DSE (calculated using $n=10$) that is an order of magnitude larger. Inspection reveals that all three have two semantic categories observed at $n=100$, and both categories are also observed within the first $10$ samples. Of these instances, all have $SE^{*}<0.05$ and one has $SE^{*}<0.005$. The impacts of this phenomenon are visible in Figures 2 and 7, where the curve for Gemma-3-12B on POTATO deviates from the pattern observed for other model-dataset pairs. If we exclude the most extreme instance by visualizing only instances with $SE^{*}\geq 0.005$, all model-dataset pairs exhibit the familiar behavior (Figure 9).
Figure 9: Illustrating underestimation in discrete semantic entropy (DSE) calculation with typical sample sizes. The ratios of semantic entropy (SE) estimators with varying sample size ($n=5,10,25,50,75,100$) to white-box SE with $n=100$ (denoted $SE^{*}$) are shown, with values below 11 suggesting underestimation (dotted grey line). The estimators displayed are the plugin estimator of canonical DSE (i.e., Equation 5, dotted green line), the “hybrid” SE estimator of Equation 10 (solid indigo line), and white-box SE (Equation 3, dotted cyan line). Results are averaged across queries within each dataset, then uniformly averaged across datasets. Log scale is used on the x-axis to highlight differences between estimators with smaller sample sizes. Instances with exceedingly small denominators (i.e., $SE^{*}<0.005$) are ignored.
Appendix E How distinct are canonical DSE and SE?
In Figures 8 and 9, we observe that SE underestimates $SE^{*}$, which is unsurprising, as it also invokes the negatively biased plugin entropy estimator (Equation 3). It is surprising, however, that SE appears to track plugin DSE so closely, since SE’s probabilities are calculated via RBMCI (Equation 4), versus plugin DSE’s empirical class frequencies. In Figure 3, SE does not appear definitively superior to plugin DSE overall on the downstream incorrectness detection task. Similarly, Farquhar et al. (2024) and Nikitin et al. (2024) report little difference in incorrectness detection performance metrics (e.g., AUROC) between SE and DSE.
In Table 4, we compare the empirical distribution of semantic class frequencies to the one induced by RBMCI, using the data from the experiment supporting Figure 3. Our metrics of comparison are their Kullback-Leibler divergence (KL) (Kullback & Leibler, 1951), total variation distance (TVD) (Jacobs, 2018), and the mean absolute error (MAE), all computed by comparing each method’s class-wise probability vectors.
Model Dataset KL TVD MAE
Gemma-2-9B HotpotQA 0.001 0.006 0.005
SQuAD 2.0 0.002 0.009 0.006
POTATO 0.004 0.021 0.014
BioASQ 0.004 0.021 0.012
Gemma-3-12B HotpotQA 0.000 0.001 0.001
SQuAD 2.0 0.000 0.002 0.002
POTATO 0.001 0.004 0.004
BioASQ 0.001 0.011 0.006
Llama-3.1-8B HotpotQA 0.009 0.025 0.018
SQuAD 2.0 0.010 0.029 0.019
POTATO 0.008 0.034 0.021
BioASQ 0.018 0.053 0.025
Mistral-v0.3-7B HotpotQA 0.001 0.009 0.007
SQuAD 2.0 0.002 0.015 0.010
POTATO 0.004 0.023 0.016
BioASQ 0.004 0.027 0.014
Phi-3.5-3.8B HotpotQA 0.003 0.013 0.010
SQuAD 2.0 0.002 0.013 0.010
POTATO 0.012 0.039 0.028
BioASQ 0.005 0.029 0.016
Table 4: Comparing semantic equivalence class probability distributions, induced by empirical frequencies (as in black-box DSE) or via Equation 4 (as in white-box SE). The metrics considered are Kullback-Leibler divergence (KL), total variation distance (TVD), and the mean absolute error between the class-wise probability vectors (MAE). For each model, we average results over the questions of: HotpotQA, the answerable questions of SQuAD 2.0, POTATO, and BioASQ. Values are rounded to four decimal places to adequately represent the smallest entries.
All three are exceedingly small across models and datasets; for instance, the KL divergence is nearly always well below $0.01$ (Table 4). We consider this finding in light of Figure 8, where the curves for DSE-based methods are visually impacted by an anomalous sampling result, but that of SE is not. We conclude that, while SE and plugin DSE may compute plugin entropy over nearly identical distributions, on average, sequence log-probabilities can help RBMCI down-weight scenarios when the sample over-represents exceptionally rare events.
Appendix F Additional incorrectness detection results